The following data relate to the Machinery account of Metlock, Inc. at December 31, 2020.

Machinery

A

B

C

D

Original cost $50,600 $56,100 $88,000 $88,000
Year purchased 2015 2016 2017 2019
Useful life 10 years 15,000 hours 15 years 10 years
Salvage value $3,410 $3,300 $5,500 $5,500
Depreciation method Sum-of-the-years'-digits Activity Straight-line Double-declining balance
Accum. depr through 2020* $34,320 $38,720 $16,500 $17,600


*In the year an asset is purchased, Metlock, Inc. does not record any depreciation expense on the asset.

In the year an asset is retired or traded in, Metlock, Inc. takes a full year’s depreciation on the asset.

The following transactions occurred during 2021.

(a) On May 5, Machine A was sold for $14,300 cash. The company’s bookkeeper recorded this retirement in the following manner in the cash receipts journal.

Cash

14,300

    Machinery (Machine A)

14,300

(b) On December 31, it was determined that Machine B had been used 2,350 hours during 2021.
(c) On December 31, before computing depreciation expense on Machine C, the management of Metlock, Inc. decided the useful life remaining from January 1, 2021, was 10 years.
(d) On December 31, it was discovered that a machine purchased in 2020 had been expensed completely in that year. This machine cost $30,800 and has a useful life of 10 years and no salvage value. Management has decided to use the double-declining-balance method for this machine, which can be referred to as “Machine E.”

I ONLY NEED PART A AND B PLEASE. :)
Answer #1

a) Accumulated Depreciation-Machinery = (50,600 - 3,410) x 5/55 = 4,290

Machinery = 50,600 - 14,300 = 36,300

Debit Credit
Depreciation Expense    4,290.00
Accumulated Depreciation    4,290.00
Accumulated Depreciation    38,610.00
Machinery    36,300.00
Gain on Disposal      2,310.00

b)

Accumulated Depreciation-Machinery = (56,100 - 3,300) x 2,350/15,000 = 8,272

Debit Credit
Depreciation Expense      8,272.00
Accumulated Depreciation      8,272.00